OpenStudy (anonymous):

Which of the following is the best linear approximation for f(x) = cos(x) near x equals pi divided by 2?

OpenStudy (anonymous):

y equals x minus pi over 2 y equals negative x plus pi over 2 y equals negative x plus pi over 2 plus 1 y equals x minus pi over 2 plus 1

OpenStudy (anonymous):

The derivative of cos(x) is -sin(x), but idk where to go from there.

OpenStudy (anonymous):

We start with this:\[ dy = \frac {dy}{dx}dx\\ \Delta y\approx \frac{dy}{dx}\Delta x \]

OpenStudy (freckles):

so we have the tangent line at x=pi/2 is \[y-f( \frac{\pi}{2})=f'( \frac{\pi}{2})(x- \frac{\pi}{2})\]

OpenStudy (freckles):

oh and that is actually what @wio wrote in fancy notation :p

OpenStudy (freckles):

\[y=f'(\frac{\pi}{2})(x-\frac{\pi}{2})+f(\frac{\pi}{2}) \\ f(x) \approx f'(\frac{\pi}{2})(x-\frac{\pi}{2})+f(\frac{\pi}{2} ) \text{ for values of } x \text{ near } \frac{\pi}{2}\]

OpenStudy (anonymous):

\[ y-y_0 \approx \frac{dy}{dx}(x-x_0) \]

OpenStudy (freckles):

so you found f'(x) already now plug pi/2 into f' to find the slope of your line

OpenStudy (anonymous):

-sin(pi/2) ?

OpenStudy (freckles):

yah do you know what that equals

OpenStudy (anonymous):

-1?

OpenStudy (freckles):

\[y=f'(\frac{\pi}{2})(x-\frac{\pi}{2})+f(\frac{\pi}{2}) \\ y=-1(x-\frac{\pi}{2})+f(\frac{\pi}{2})\]

OpenStudy (freckles):

cool

OpenStudy (freckles):

now one step really left find f(pi/2)

OpenStudy (freckles):

plug pi/2 into your original function that is

OpenStudy (anonymous):

-x+pi/2 is the answer? correct?

OpenStudy (freckles):

:)

OpenStudy (anonymous):

thank you i made that way harder than it was lol

OpenStudy (freckles):

no problem